# Investigating and Defending Shortcut Learning in Personalized Diffusion Models

### v3

@@ Appendix B More Experimental Results
Different Negative Tokens. To investigate the effect of using different negative tokens, we additionally present results in Table 5 and Figure 7. As we can see, setting the right negative tokens is crucial for the performance. In our four tested prompts, we found that “UNKNOWN artifact masked” yielded the best performance in both IMS and Q scores. Future works can be conducted using automatic negative prompt searching. Another direction is to study visualization of the learned pattern for each negative prompt setting for a deeper understanding of personalized diffusion models.
Figure 7: Scores vs Prompt Index.
Table 5: Evaluation on our method with different negative tokens.
IMS Q Overall Score #Prompt Index Negative Tokens
0.305 0.655 0.480 0 UNKNOWN degraded quality
0.367 0.682 0.525 1 UNKNOWN artifact masked
0.148 0.550 0.349 2 UNKNOWN noise corruption
-0.134 0.426 0.146 3 BHI noise corruption
Figure 8: Evaluation of trained models on perturbed and purified images across perturbation radii.
Figure 9: More visualization of perturbed and purified images.
Appendix C Implementation of Baselines
C.1 Purification Methods
@@
We implement two classes of purification approaches; the first ones are model-free and operate with certain image processing algorithms, such as Gaussian Filtering, total variation minimization (TVM), and JPEG compression. Despite the simplicity, researchers found that these approaches can achieve non-trivial defense performance against adversarial attacks [7], availability attacks [8, 4], and more general data poisoning attacks [53]. Another line of approach is based on powerful diffusion probabilistic models, which have a strong ability to model real-world data distribution and also show potential in being leveraged for zero-shot purifiers [54, 13, 49, 12]. We include a wide range of SoTA diffusion-based purification approaches that are designed for the protective perturbation specifically, including GrIDPure [15], IMPRESS [12], or those are proposed for more general adversarial perturbation [14, 49], including DiffPure [14] (with pixel-space diffusion models or latent-space diffusion models), DDS-based purification (DDSPure) [49].
@@
1. Gaussian Filtering. Gaussian Filtering is a well-known image-processing technique used to reduce image noise and detail by applying a Gaussian kernel. The high-frequent part in adversarial perturbation can be smoothed after filtering. We set the kernel size as 5 following [4].
@@
2. Total Variation Minimization (TVM) [55] The main idea of TVM is to conduct image reconstruction based on the observation that the benign images should have low total variation. We implemented the TVM defense in the following steps: we first resized the instance image to 64x64 pixels, applied a random dropout mask with a 2% pixel dropout rate, and solved a TVM optimization problem. The optimization aims to reconstruct the image by minimizing the difference between the original and reconstructed images while enforcing smoothness through the total variation term: $\min_{Z}||(1-X)\odot(Z-x)||_{2}+\lambda_{TV}TV_{2}(Z).$ After optimization, the reconstructed image is reshaped back to $6464x$ 64 and then upsampled to $512512x$ 512 through two SR steps with a middle resizing process.
@@
3. JPEG Compression. It involves transforming an image into a format that uses less storage space and reduces the image file’s size. We set the JPEG quality to 75 following [8].
@@
4. DiffPure [14]. Diffusion Purification (DiffPure) first diffuses the adversarial example with a small amount of noise given a pre-defined timestep $t$ following a forward diffusion process, where the adversarial noise is smoothed and then recovers the clean image through the reverse generative process. Depending on the type of diffusion model used, this simple yet effective approach can be adapted into two versions: PDM-based DiffPure and LDM-based DiffPure. In our implementation, we term the PDM-based DiffPure as PixelDiffPure for short and leverage 256x256_diffusion_uncond pre-trained on ImageNet released in the guided-diffusion following common practice. For the LDM-based DiffPure, we term it as LatentDiffPure since the diffusion process is conducted in latent space and leverage Stable Diffusion v1-4 [50] for its superior performance. Since the SD model has the ability to input additional text prompts during the purification process, we investigate two variants with and without the usage of purified text prompting. For LatentDiffPure-$\emptyset$, we set the text to null, while for LatentDiffPure, we set it as ‘‘a photo of [class_name], high quality, highres’’.
@@
5. DDSPure [49]. Similar to DiffPure [14], the main idea behind Diffusion Denoised Smoothing (DDS) is to find an optimal timestamp that can maximally remove the adversarial perturbation via the SDEdit process [31]. Given smoothing noise level $\delta$, the optimal timestamp $t^{*}$ is computed via, $t^{*}=\frac{1-\bar{\alpha_{t}}}{\bar{\alpha_{t}}}=\sigma^{2}.$ Following common practice, we leverage the pre-trained diffusion model on ImageNet released in the guided-diffusion. Specifically, the 256x256_diffusion_uncond is used as a denoiser. To resolve the size mismatch, we resize the images to fit the model input and resize the image size back after purification. And we clip $t^{*}$ when it falls outside the sampling step range of $[0,1000]$.
@@
4. GrIDPure [15]. GrIDPure notices that for purification in defending protective perturbation, conducting iterative DiffPure with small steps can outperform one-shot DiffPure with larger steps. Furthermore, it suppresses the generative nature during diffusion purification by additionally splitting the image into multiple small grids that are separately processed with a final merging process. This allows the model to focus more on purifying those perturbed textures and curves in the image without mistakenly affecting the overall structure, thus preserving the faithfulness of purification.
@@
These steps constitute one iteration, and the algorithm is repeated for a total of $10$ iterations.
6. IMPRESS [12] The key idea of IMPRESS is to conduct purification that ensures latent consistency with visual similarity constraints: (1) the purified image should be visually similar to the perturbed image, and (2) the purified image should be consistent upon a LDM-based reconstruction. To quantify the similarity condition, IMPRESS uses the LPIPS metric [33], which measures the human-perceived image distortion between the purified image $\mathbf{x}_{\text{pur}}$ and the perturbed image $\mathbf{x}_{\text{ptb}}$. The loss is defined as $\max(\text{LPIPS}(\mathbf{x}_{\text{pur}},\mathbf{x}_{\text{ptb}})-\Delta_{L},0)$, where $\Delta_{L}$ is the perceptual perturbation budget. For the consistency condition, IMPRESS simplifies the loss by removing the diffusion process and defines it as $||\mathbf{x}_{\text{pur}}-\mathcal{D}(\mathcal{E}(\mathbf{x}_{\text{pur}}))||^{2}_{2}$, where $\mathcal{E}$ and $\mathcal{D}$ are the image encoder and decoder in the LDM, respectively. The final optimization problem combines the two losses:
@@
where $\alpha$ is a hyperparameter to balance the two losses, which is set as $0.1$. The optimization is solved with PGD [56] with Adam optimizer with lr of 0.001, and the total iteration is set as 3000.
C.2 Protective Perturbation Methods

### v4

@@ B.2 Learned Concepts Visualization
To investigate the effect of using our CDL with different noise tokens, we additionally present results in Tab. 5. As we can see, setting the right noise tokens is crucial for the performance. Among the 15 noise tokens we tested, we found that “t@j noisy pattern” yielded the best overall performance under our setting. Future works can be conducted using automatic noise prompt searching. Another direction is to study visualization of the learned pattern for each noise prompt setting for a deeper understanding of the underlying concept learning process.
Table 5: Performance comparison of models trained with our Contrastive Decoupling Learning (CDL) using various noise tokens. Results are shown for seven evaluation metrics across different noise token choices. Higher scores indicate better performance. The experiment uses a single random instance from the VGGFace2 dataset, protected by MetaCloak. IMS and Q are our main metrics, while $\text{IMS}_{\text{VGG}}$, $\text{IMS}_{\text{IP}}$, LIQE, and CLIP-IQAC provide additional insights into model performance.
Noise Tokens $\mathcal{V}_{N}*$ IMS $\uparrow$ Q $\uparrow$ Avg. $\uparrow$ $\text{IMS}_{\text{VGG}}$uparrow $\text{IMS}_{\text{IP}}$uparrow LIQE $\uparrow$ CLIP-IQAC $\uparrow$
t@j noisy pattern -0.226 0.156 -0.035 -0.265 -0.209 3.313 0.460
xjy image imperfection -0.331 -0.130 -0.230 -0.511 -0.253 2.740 0.324
xjy visual interference -0.539 0.014 -0.263 -0.513 -0.550 3.028 0.130
xjy visual distortion -0.336 -0.204 -0.270 -0.284 -0.357 2.591 0.149
xjy image artifact -0.159 -0.445 -0.302 -0.423 -0.045 2.110 0.277
xjy digital glitch -0.294 -0.378 -0.336 -0.448 -0.227 2.243 0.328
UNKNOWN face degradation -0.197 -0.476 -0.337 -0.296 -0.155 2.047 0.520
xjy image disturbance -0.328 -0.366 -0.347 -0.424 -0.287 2.268 0.225
xjy image corruption -0.449 -0.248 -0.349 -0.477 -0.437 2.504 0.104
xjy image degradation -0.345 -0.410 -0.377 -0.286 -0.370 2.181 0.110
UNKNOWN noisy pattern -0.431 -0.419 -0.425 -0.244 -0.512 2.161 -0.020
xjy visual anomaly -0.389 -0.534 -0.461 -0.453 -0.361 1.932 0.126
XX noisy artifact -0.324 -0.626 -0.475 -0.143 -0.401 1.749 -0.096
xjy visual noise -0.578 -0.475 -0.526 -0.288 -0.702 2.050 -0.060
bhi noisy perturbation -0.494 -0.727 -0.610 -0.369 -0.547 1.546 -0.242
Table 6: The full results with standard deviations of different purification methods under different protective perturbations. The best performances are in bold, and second runners are shaded in gray. * denotes improvement that passes the Wilcoxon signed-rank significance test with $p\leq 0.01$.
Methods FSMG ASPL EASPL MetaCloak AdvDM PhotoGuard Glaze
IMS $\uparrow$ Q $\uparrow$ IMS Q IMS Q IMS Q IMS Q IMS Q IMS Q
Clean -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08 -0.13 $\pm$ 0.04 0.15 $\pm$ 0.08
Perturbed -0.43 $\pm$ 0.54 -0.54 $\pm$ 0.25 -0.67 $\pm$ 0.46 -0.52 $\pm$ 0.41 -0.62 $\pm$ 0.46 -0.50 $\pm$ 0.40 -0.35 $\pm$ 0.58 -0.53 $\pm$ 0.28 -0.27 $\pm$ 0.54 -0.36 $\pm$ 0.30 -0.18 $\pm$ 0.54 -0.24 $\pm$ 0.27 -0.28 $\pm$ 0.59 -0.28 $\pm$ 0.33
Gaussian F. -0.19 $\pm$ 0.57 -0.55 $\pm$ 0.29 -0.20 $\pm$ 0.56 -0.57 $\pm$ 0.20 -0.17 $\pm$ 0.56 -0.58 $\pm$ 0.23 -0.07$\pm$ 0.54 -0.63 $\pm$ 0.15 -0.11 $\pm$ 0.54 -0.57 $\pm$ 0.24 -0.23 $\pm$ 0.56 -0.53 $\pm$ 0.26 -0.18 $\pm$ 0.53 -0.54 $\pm$ 0.25
JPEG -0.15 $\pm$ 0.60 -0.41 $\pm$ 0.33 -0.21 $\pm$ 0.62 -0.52 $\pm$ 0.25 -0.27 $\pm$ 0.62 -0.50 $\pm$ 0.24 -0.34 $\pm$ 0.63 -0.38 $\pm$ 0.41 -0.15 $\pm$ 0.62 -0.02$\pm$ 0.39 -0.13 $\pm$ 0.62 0.07$\pm$ 0.36 -0.19 $\pm$ 0.57 -0.03$\pm$ 0.48
TVM -0.12 $\pm$ 0.48 -0.65 $\pm$ 0.21 -0.16 $\pm$ 0.53 -0.66 $\pm$ 0.23 -0.10 $\pm$ 0.49 -0.67 $\pm$ 0.20 -0.11 $\pm$ 0.49 -0.69 $\pm$ 0.12 -0.12 $\pm$ 0.50 -0.65 $\pm$ 0.22 -0.15 $\pm$ 0.54 -0.64 $\pm$ 0.20 -0.11 $\pm$ 0.48 -0.66 $\pm$ 0.20
PixelDiffPure -0.41 $\pm$ 0.45 -0.57 $\pm$ 0.17 -0.43 $\pm$ 0.47 -0.54 $\pm$ 0.21 -0.57 $\pm$ 0.49 -0.61 $\pm$ 0.16 -0.28 $\pm$ 0.44 -0.58 $\pm$ 0.22 -0.40 $\pm$ 0.51 -0.55 $\pm$ 0.21 -0.25 $\pm$ 0.51 -0.55 $\pm$ 0.15 -0.41 $\pm$ 0.44 -0.59 $\pm$ 0.17
L.DiffPure-$\emptyset$ -0.07$\pm$ 0.50 -0.47 $\pm$ 0.29 -0.36 $\pm$ 0.49 -0.59 $\pm$ 0.21 -0.22 $\pm$ 0.58 -0.49 $\pm$ 0.26 -0.52 $\pm$ 0.46 -0.43 $\pm$ 0.29 -0.55 $\pm$ 0.45 -0.24 $\pm$ 0.38 -0.12 $\pm$ 0.48 -0.40 $\pm$ 0.28 -0.38 $\pm$ 0.45 -0.42 $\pm$ 0.27
L.DiffPure -0.25 $\pm$ 0.48 -0.45 $\pm$ 0.30 -0.31 $\pm$ 0.51 -0.61 $\pm$ 0.20 -0.30 $\pm$ 0.54 -0.46 $\pm$ 0.32 -0.31 $\pm$ 0.46 -0.51 $\pm$ 0.22 -0.57 $\pm$ 0.43 -0.30 $\pm$ 0.34 -0.25 $\pm$ 0.48 -0.47 $\pm$ 0.21 -0.41 $\pm$ 0.46 -0.47 $\pm$ 0.27
DDSPure -0.15 $\pm$ 0.61 -0.34 $\pm$ 0.23 -0.05$\pm$ 0.59 -0.38 $\pm$ 0.19 -0.08$\pm$ 0.54 -0.39 $\pm$ 0.20 -0.16 $\pm$ 0.59 -0.49 $\pm$ 0.23 -0.19 $\pm$ 0.59 -0.43 $\pm$ 0.23 -0.12 $\pm$ 0.59 -0.37 $\pm$ 0.21 -0.22 $\pm$ 0.55 -0.41 $\pm$ 0.24
GrIDPure -0.10 $\pm$ 0.59 -0.20$\pm$ 0.23 -0.21 $\pm$ 0.58 -0.16$\pm$ 0.23 -0.13 $\pm$ 0.55 -0.25$\pm$ 0.25 -0.23 $\pm$ 0.52 -0.25$\pm$ 0.25 -0.09$\pm$ 0.54 -0.18 $\pm$ 0.26 -0.03$\pm$ 0.59 -0.22 $\pm$ 0.26 -0.24 $\pm$ 0.56 -0.13 $\pm$ 0.28
IMPRESS -0.15 $\pm$ 0.58 -0.53 $\pm$ 0.24 -0.16 $\pm$ 0.60 -0.49 $\pm$ 0.31 -0.29 $\pm$ 0.60 -0.64 $\pm$ 0.13 -0.34 $\pm$ 0.58 -0.29 $\pm$ 0.30 -0.34 $\pm$ 0.56 -0.34 $\pm$ 0.31 -0.16 $\pm$ 0.58 -0.21 $\pm$ 0.28 -0.10$\pm$ 0.59 -0.43 $\pm$ 0.25
Ours 0.23 $\pm$ 0.47 0.65 $\pm$ 0.21 0.09 $\pm$ 0.48 0.62 $\pm$ 0.15 0.09 $\pm$ 0.49 0.63 $\pm$ 0.19 0.38 $\pm$ 0.38 0.58 $\pm$ 0.27 0.29 $\pm$ 0.44 0.67 $\pm$ 0.20 0.24 $\pm$ 0.49 0.63 $\pm$ 0.19 0.31 $\pm$ 0.43 0.66 $\pm$ 0.25
B.4 More Results on Purification Faithfulness
In addition to the LPIPS used in the paper, we provide purification results using other similarity metrics, including Structural Similarity (SSIM), Multi-Scale Structural Similarity (MS-SSIM), and Peak Signal-to-Noise Ratio (PSNR). The results are presented in Tab.  7, which demonstrate that our purification variants are consistently superior to previous state-of-the-art purification approaches.
Table 7: Purification faithfulness under various similarity metrics.
Settings LPIPS $\downarrow$ SSIM $\uparrow$ MS_SSIM $\uparrow$ PSNR $\uparrow$ Avg(IMS,Q) $\uparrow$
IMPRESS 0.451 0.761 0.903 49.294 -0.63
DDSPure 0.384 0.805 0.873 46.948 -0.65
GrIDPure 0.429 0.754 0.794 41.976 -0.48
L.DiffPure-$\emptyset$ 0.450 0.676 0.732 43.551 -0.82
Code + SR 0.271 0.824 0.925 49.937 0.385
Code + SR 0.231 0.891 0.952 52.49 0.339
Code + SR 0.270 0.790 0.923 49.591 0.225
B.5 Limitations Discussion
While our proposed defense framework demonstrates significant improvements over existing methods in enhancing the robustness of PDMs, there are certain areas that could be further explored. Our experiments are primarily conducted on the facial dataset VGGFace2. Although we have preliminary purification results indicating the applicability of our approach to other domains like artwork images from WikiArt, we have not extensively tested our method across a wide variety of protection techniques. Future work could investigate the generalizability of our method to different types of images and subjects to further validate its effectiveness. Additionally, the integration of data purification and contrastive decoupling learning introduces some additional computational steps during the training process. This may slightly increase the training time compared to standard training procedures. However, we believe that this is a reasonable trade-off given the substantial benefits in terms of robustness and generation quality that our method provides. While our framework shows robustness against certain adaptive perturbations, it is always possible that more sophisticated adversarial strategies could be developed. Ongoing research and adaptive strategies may be necessary to continually strengthen the model’s resilience against evolving threats.
Appendix C Causal Analysis of Learning Personalized Diffusion Models on Perturbed Data
C.1 Construction of the Causal Graph when Learning PDMs on Perturbed Data
To understand how adversarial perturbations lead to shortcut learning in PDMs, we construct a Structural Causal Model (SCM) that captures the causal relationships between the variables involved in the fine-tuning process. The variables in our SCM are defined as follows: $\mathbf{X}$ represents the original clean images representing the true concept; $\bm{\delta}$ denotes the adversarial perturbations added to the images; $\mathbf{X}^{\prime}=\mathbf{X}+\bm{\delta}$ are the perturbed images used for fine-tuning; $\bm{c}$ represents class-specific text prompts without the unique identifier (e.g., “a photo of a person”); $\mathcal{V}^{*}$ is the unique identifier token used in personalized prompts (e.g., “sks”); $\bm{c}^{\mathcal{V}^{*}}=\bm{c}\oplus\mathcal{V}^{*}$ denotes the personalized text prompts combining $\bm{c}$ and $\mathcal{V}^{*}$; $\theta$ represents the model parameters being fine-tuned, and $\mathbf{Y}$ denotes the model outputs (generated images). The structural equations governing the relationships are: The structural equations governing the relationships in our SCM are as follows: (1) Perturbed Images: $\mathbf{X}^{\prime}=\mathbf{X}+\bm{\delta}$, where $\mathbf{X}^{\prime}$ represents the perturbed images, $\mathbf{X}$ the original clean images, and $\bm{\delta}$ the adversarial perturbations. (2) Model Fine-tuning: $\theta=f_{\theta}(\theta_{0},\mathbf{X}^{\prime},\bm{c}^{\mathcal{V}^{*}})$, where $\theta$ represents the fine-tuned model parameters, $\theta_{0}$ the initial model parameters, and $\bm{c}^{\mathcal{V}^{*}}$ the personalized text prompts. (3) Model Output Generation: $\mathbf{Y}=g(\theta,\bm{c}^{\mathcal{V}^{*}})$, where $\mathbf{Y}$ denotes the generated images.
Based on these equations, we construct the causal graph (model $\theta$ and output $\mathbf{Y}$ omitted for clarity) shown in Fig. 2. In the graph, the adversarial perturbations $\bm{\delta}$ directly affect the perturbed images $\mathbf{X}^{\prime}$. The perturbed images $\mathbf{X}^{\prime}$ and personalized prompts $\bm{c}^{\mathcal{V}^{*}}$ influence the fine-tuned model parameters $\theta$. The fine-tuned model parameters $\theta$ and the personalized prompts $\bm{c}^{\mathcal{V}^{*}}$ influence the generated outputs $\mathbf{Y}$. Spurious correlations (shortcut paths) are introduced from $\bm{\delta}$ to $\theta$ and subsequently to $\mathbf{Y}$, leading the model to incorrectly associate the adversarial noise with the unique identifier $\mathcal{V}^{*}$.
C.2 Connection of Our Design to Causal Intervention
Our defense strategies can be interpreted as interventions that modify the causal graph to weaken or eliminate the undesired influence of adversarial perturbations $\bm{\delta}$ on the model outputs $\mathbf{Y}$.
Input Purification aims to mitigate the effect of adversarial perturbations, but it’s important to note that this process is not perfect. In the context of our causal model, we can represent this imperfect purification as:
$\mathbf{X}^{\prime}=\mathbf{X}+\bm{\delta}\rightarrow\mathbf{X}^{\prime}=\mathbf{X}+\bm{\delta}_{r},$ (7)
where $\bm{\delta}_{r}$ represents the residual perturbations after purification, with $\|\bm{\delta}_{r}\|\ll\|\bm{\delta}\|$. This intervention partially weakens the path from $\bm{\delta}$ to $\mathbf{X}^{\prime}$, and consequently to $\theta$ and $\mathbf{Y}$. The distribution of the model outputs given this imperfect purification can be expressed as:
$P(\mathbf{Y}\mid\text{do}(\bm{\delta}=\bm{\delta}_{r}),\mathbf{X},\bm{c}^{\mathcal{V}^{*}}).$ (8)
While input purification reduces the influence of adversarial perturbations on the fine-tuning process and subsequent image generation, it does not completely eliminate the shortcut learning problem. This limitation motivates the need for additional strategies to mitigate shortcut learning further.
Contrastive Decoupling Learning (CDL) intervenes on the $\bm{\delta}$$\mathcal{V}^{*}$ relationship by introducing noise tokens $\mathcal{V}_{N}^{*}$, thus weakening $\bm{\delta}$’s indirect effect on $\theta$. By augmenting the instance prompts to include the noise token (e.g., “a photo of $\mathcal{V}^{*}$ with $\mathcal{V}_{N}^{*}$ noisy pattern”) and the class prompts to exclude it (e.g., “a photo of a person without $\mathcal{V}_{N}^{*}$ noisy pattern”), CDL encourages the model to disentangle the learning of the personalized concept from the noise patterns. Furthermore, when we apply the quality-enhanced sampling with CFG, it modifies the generation process by incorporating negative prompts during inference, thereby adjusting the output equation to:
$\mathbf{Y}=g^{\prime}(\theta,\bm{c}^{\mathcal{V}^{*}},\bm{c}_{\text{neg}}),$ (9)
where $g^{\prime}$ is the modified generation function and $\bm{c}_{\text{neg}}$ are negative prompts (e.g., “noisy, abstract, pattern, low quality”). This step acts as an intervention on the generation mechanism, reducing the influence of any residual associations between $\bm{\delta}$ and the outputs. Although it is more of a practical adjustment than a formal causal intervention, it helps steer the model toward generating high-quality images that reflect the true personalized concept.
By combining these strategies, we provide a comprehensive approach to mitigate shortcut learning in PDMs. Input purification directly removes the influence of adversarial perturbations, CDL reduces spurious associations during training, and quality-enhanced sampling improves the quality of generated outputs by discouraging undesired artifacts.
Appendix D Implementation of Baselines
D.1 Purification Methods
@@
We implement two classes of purification approaches; the first ones are model-free and operate with certain image processing algorithms, such as Gaussian Filtering, total variation minimization (TVM), and JPEG compression. Despite the simplicity, researchers found that these approaches can achieve non-trivial defense performance against adversarial attacks (Liang et al., 2023), availability attacks (Liu et al., 2024; Van Le et al., 2023), and more general data poisoning attacks (Huang et al., 2020). Another line of approach is based on powerful diffusion probabilistic models, which have a strong ability to model real-world data distribution and also show potential in being leveraged for zero-shot purifiers Shi et al. (2024); Zhao et al. (2024); Carlini et al. (2022); Cao et al. (2024). We include a wide range of SoTA diffusion-based purification approaches that are designed for the protective perturbation specifically, including GrIDPure (Zheng et al., 2023), IMPRESS (Cao et al., 2024), or those are proposed for more general adversarial perturbation (Nie et al., 2022; Carlini et al., 2022), including DiffPure (Nie et al., 2022) (with pixel-space diffusion models or latent-space diffusion models), DDS-based purification (DDSPure) (Carlini et al., 2022; Hu et al., ).
@@
1. Gaussian Filtering. Gaussian Filtering is a well-known image-processing technique used to reduce image noise and detail by applying a Gaussian kernel. The high-frequent part in adversarial perturbation can be smoothed after filtering. The kernel size is set as 5 following (Van Le et al., 2023).
@@
2. Total Variation Minimization (TVM) (Wang et al., 2020) The main idea of TVM is to conduct image reconstruction based on the observation that the benign images should have low total variation. We implemented the TVM defense in the following steps: we first resized the instance image to $64^{2}$ pixels, applied a random dropout mask with a 2 After optimization, the reconstructed image is reshaped back to $64^{2}$ and then upsampled to $512^{2}$ through two SR steps with a middle resizing process.
@@
3. JPEG Compression. It involves transforming an image into a format that uses less storage space and reduces the image file’s size. We set the JPEG quality to 75 following (Liu et al., 2024).
@@
4. DiffPure (Nie et al., 2022). Diffusion Purification (DiffPure) first diffuses the adversarial example with a small amount of noise given a pre-defined timestep $t$ following a forward diffusion process, where the adversarial noise is smoothed and then recovers the clean image through the reverse generative process. Depending on the type of diffusion model used, this simple yet effective approach can be adapted into two versions: PDM-based DiffPure and LDM-based DiffPure. In our implementation, we term the PDM-based DiffPure as PixelDiffPure for short and leverage 256x256_diffusion_uncond pre-trained on ImageNet released in the guided-diffusion following common practice. For the LDM-based DiffPure, we term it as LatentDiffPure since the diffusion process is conducted in latent space and leverage Stable Diffusion v1-4 (Rombach et al., 2022) for its superior performance. Since the SD model has the ability to input additional text prompts during the purification process, we investigate two variants with and without the usage of purified text prompting. For LatentDiffPure-$\emptyset$, we set the text to null, while for LatentDiffPure, we set it as ‘‘a photo of [class_name], high quality, highres’’.
@@
5. DDSPure (Carlini et al., 2022). Similar to DiffPure (Nie et al., 2022), the main idea behind Diffusion Denoised Smoothing (DDS) is to find an optimal timestamp that can maximally remove the adversarial perturbation via the SDEdit process (Meng et al., 2021). Given smoothing noise level $\delta$, the optimal timestamp $t^{*}$ is computed via, $t^{*}=\frac{1-\bar{\alpha_{t}}}{\bar{\alpha_{t}}}=\sigma^{2}.$ Following common practice, we leverage the pretrained diffusion model on ImageNet released in the guided-diffusion. Specifically, the 256x256_diffusion_uncond is used as a denoiser. To resolve the size mismatch, we resize the images to fit the model input and resize the image size back after purification. And we clip $t^{*}$ when it falls outside the sampling step range of $[0,1000]$.
@@
4. GrIDPure (Zheng et al., 2023). GrIDPure notices that for purification in defending protective perturbation, conducting iterative DiffPure with small steps can outperform one-shot DiffPure with larger steps. Furthermore, it suppresses the generative nature during diffusion purification by additionally splitting the image into multiple small grids that are separately processed with a final merging process. This allows the model to focus more on purifying those perturbed textures and curves in the image without mistakenly affecting the overall structure, thus preserving the faithfulness of purification.
@@
These steps constitute one iteration, and the algorithm is repeated for a total of $10$ iterations. We implement the GrIDPure algorithm following their official implementation.
6. IMPRESS (Cao et al., 2024) The key idea of IMPRESS is to conduct purification that ensures latent consistency with visual similarity constraints: (1) the purified image should be visually similar to the perturbed image, and (2) the purified image should be consistent upon an LDM-based reconstruction. To quantify the similarity condition, IMPRESS uses the LPIPS metric (Zhang et al., 2018), which measures the human-perceived image distortion between the purified image $\mathbf{x}_{\text{pur}}$ and the perturbed image $\mathbf{x}_{\text{ptb}}$. The loss is defined as $\max(\text{LPIPS}(\mathbf{x}_{\text{pur}},\mathbf{x}_{\text{ptb}})-\Delta_{L},0)$, where $\Delta_{L}$ is the perceptual perturbation budget. For the consistency condition, IMPRESS simplifies the loss by removing the diffusion process and defines it as $||\mathbf{x}_{\text{pur}}-\mathcal{D}(\mathcal{E}(\mathbf{x}_{\text{pur}}))||^{2}_{2}$, where $\mathcal{E}$ and $\mathcal{D}$ are the image encoder and decoder in the LDM, respectively. The final optimization problem combines the two losses:
@@
where $\alpha$ is a hyperparameter to balance the two losses, which is set as $0.1$. The optimization is solved with PGD (Madry et al., 2018) with Adam optimizer with lr of 0.001, and the total iteration is set as 3000.
D.2 Protective Perturbation Methods